OpenStudy (anonymous):

Can someone please help me with this one I am not good with graphs. IS THE LAST ANSWER CORRECT? Select the system of equations that corresponds to the given graph. Choices: 4x + y = 6 6x + 3y = 6 −3x + y = 2 4x + 2y = 4 2x + y = 4 6x + 3y = 12 4x + 2y = 6 6x + 3y = 12

OpenStudy (anonymous):

Here is the graph

OpenStudy (anonymous):

why the graph has only 1 line??? if it is a sistem of 2 equations it must have 2 lines

OpenStudy (anonymous):

u cn see that graph is passing by (0,4) and (2,0) and both these points is satisfied by option c) 2x + y = 4 6x + 3y = 12

OpenStudy (anonymous):

Oh ok, thank you and yes it only has one line.

OpenStudy (anonymous):

at least that the equations are the same for example 2x + y = 4 6x + 3y = 12 are the same line

OpenStudy (anonymous):

Ok I appreciate it. Sorry internet disconnected.